OpenStudy (anonymous):

(Geometry) Write the equation of the circle in standard form with center and radius given. Center (5, -3), radius = 1

OpenStudy (zzr0ck3r):

(5-x)^2 + (-3-y)^2 = 1

OpenStudy (zzr0ck3r):

do you see the pattern?

OpenStudy (anonymous):

Now I do.

OpenStudy (zzr0ck3r):

if the radius was 2 it would have been = 2^2 is always = to radius^2 but 1^2 = 1 ....

OpenStudy (anonymous):

Genius.

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

Lol, appreciate it sir.

OpenStudy (zzr0ck3r):

np